# Stochastic Semi-Gradient Descent for Learning Mean Field Games with Population-Aware Function Approximation

### v1

@@ A.2.1 Comparison of PA-LFA and grid discretization
To demonstrate the effectiveness of PA-LFA, we compare it with grid discretization. The reference MFE is calculated using a grid discretization with a granularity of $1/50$. We only apply LFA to the population measure estimate, as this is our main focus. The measure basis is chosen as
$\psi_{i}(s)=\mathcal{N}(0)-\mathcal{N}(\tan((s-s_{i})\pi)),$ (25)
@@
where $\mathcal{N}$ is the normal distribution, and $s_{i}$ is the center of the basis function. We run SemiSGD $d_{2}$ states (grid discretization) and $d_{2}$ basis functions (PA-LFA) for $T=10^{4}$ steps. Other parameters are the same as the general setup. The final MSE (accuracy) comparison is reported in Table 2. The MSE curves and learned population distributions are shown in Figure 6 and Figure 5, respectively. As we can see, when $d_{2}$ is small, PA-LFA generalizes better than grid discretization, as the representation capacity of $\operatorname{span}\psi$ is larger than the grid discretization. However, when $d_{2}$ is large, i.e., the discretization is fine enough, grid discretization performs better than PA-LFA. This is because the PA-LFA approximation error is larger than the discretization error when the discretization is fine enough.
@@
A.3 Routing game on a network.
We set $c_{1}=10^{5}$ and $c_{2}=10$. The other algorithmic parameters are chosen as follows: the discount factor $\gamma=0.8$, the initial state is uniformly sampled, the initial value function is set as all-zero, the initial population is randomly generated.
Similarly, ER is implemented with a low inverse temperature of $10^{-3}$, while all other methods use an inverse temperature of $10^{2}$. The performance comparison is reported in Figure 7. Notably, all methods except FPI with ER have an oscillatory behavior in the exploitability around the same value. This is due to the highly non-smooth nature of the underlying environment and the choice of near-greedy policies. Applying ER to all methods results in similar performance, as shown in Figure 8.
(a) MSE.
(b) Exploitablity.
(c) Learned population distribution.
Figure 7: Network routing.
(a) MSE under ER.
(b) Exploitablity under ER.
Figure 8: Network routing with regularization.
Appendix B Notation
@@
Backtracking is an analysis technique introduced to tackle rapidly changing Markov chains [40, 39, 38]. It considers a virtual stationary Markov chain by backtracking a period $\tau$, fixing the parameter $\xi_{t-\tau}$, and then sampling the Markovian observations with the fixed parameter. By the ergodicity of stationary Markov chains (Assumption 3), the virtual trajectory rapidly converges to the stationary distribution induced by $\xi_{t-\tau}$. We denote $\widetilde{O}_{t}$ as the virtual observation tuple on this virtual trajectory at time $t$. When we consider the semi-gradients on this virtual trajectory, we write out its dependence on $\widetilde{O}_{t}$ explicitly:
@@
$\bar{\mathfrak{g}}_{\circ}=\bar{\mathfrak{g}}_{\xi_{\circ}},\quad\bar{A}_{\circ}=\bar{A}_{\xi_{\circ}},\quad\bar{b}_{\circ}=\bar{b}_{\xi_{\circ}},\quad\bar{\psi}_{\circ}=\bar{\psi}_{\xi_{\circ}}.$ (35)
@@
Appendix C Proof of Proposition 1
@@
$\mu^{\pi}_{M}$ $(s^{\prime})=\int_{\mathcal{S}\times\mathcal{A}}P(s^{\prime}\,|\,s,a,M)\pi(a\,|\,s)\mu^{\pi}_{M}(s)\mathrm{d}s\mathrm{d}a=\int_{\mathcal{S}\times\mathcal{A}}\left<\phi(s,a),\Omega_{M}\psi(s^{\prime})\right>\pi(a\,|\,s)\mu^{\pi}_{M}(s)\mathrm{d}s\mathrm{d}a$ (37)
$=$ $\int_{\mathcal{S}\times\mathcal{A}}\left<\psi(s^{\prime}),\Omega_{M}^{T}\phi(s,a)\right>\pi(a\,|\,s)\mu^{\pi}_{M}(s)\mathrm{d}s\mathrm{d}a=\left<\psi(s^{\prime}),\underbrace{\int_{\mathcal{S}\times\mathcal{A}}\Omega_{M}^{T}\phi(s,a)\pi(a\,|\,s)\mu_{M}^{\pi}(s)\mathrm{d}s\mathrm{d}a}_{\eta}\right>.$
Appendix D Projected MFE and stationary point
We call a value function-population measure pair a stationary point if the mean-path semi-gradient evaluated at this point is zero. Sections 5 and F show that SemiSGD converges to a stationary point. Ideally, we want the stationary point to be an MFE (Definition 1). Propositions 1 and C show that for linear MFGs, images of $\mathcal{T}$ and $\mathcal{P}$ are within the linear span of $\phi$ and $\psi$, indicating the linear structure of the MFE. However, this does not hold for general (non-linear) MFGs, hinting the discrepancy between the stationary point and the MFE for non-linear MFGs.
@@ Appendix D Projected MFE and stationary point
where $\Pi_{\phi}$ and $\Pi_{\psi}$ are orthogonal projection operators onto the linear span of $\phi$ and $\psi$, respectively.
@@
where $\widehat{G}_{\xi}$ is the Gram matrix of $\phi$ w.r.t. the inner product $\langle\cdot,\cdot\rangle_{\mu_{\xi}}$. Note that $\widehat{G}_{\xi}$ is different from the TD operator ${G}_{\phi}$ or $G_{\xi}$ defined in Appendix B, which is only Gram-like.
@@
Appendix E Preliminary lemmas
@@
We first prove the last inequality in the lemma. By Mitrophanov [28, Corollary 3.1], we have
@@
Appendix F Sample complexity analysis
F.1 Key lemmas
@@
$w_{\min}(\bar{G}_{\xi_{*}})=\sup_{\|\theta\|=1}\mathbb{E}_{\xi_{*}}\left[\theta^{T}\phi(s,a)(\phi(s,a)-\gamma\phi(s^{\prime},a^{\prime}))^{T}\theta\right]\eqqcolon\sup_{\|\theta\|=1}\mathbb{E}_{\xi_{*}}\left[u^{2}-\gamma uu^{\prime}\right],$ (122)
@@
$w_{\min}(\bar{G}_{\xi_{*}})\geq(1-\gamma)\sup_{\|\theta\|=1}\mathbb{E}_{\xi_{*}}[u^{2}]=(1-\gamma)w_{\min}\left(\mathbb{E}_{\xi_{*}}[\phi\phi^{T}]\right)=(1-\gamma)w_{\min}\left(\hat{G}_{\xi_{*}}\right)=(1-\gamma)\lambda_{\min}\left(\hat{G}_{\xi_{*}}\right),$ (124)
where $\hat{G}_{\xi_{*}}$ is the Gram matrix of the feature map $\phi$ under the steady distribution induced by $\xi_{\xi_{*}}$, and the last equality uses the property of normal matrices. Plugging the above derivation into Equation 121 gives
@@
It is clear that Equations 125 and 116 are controlled by both $\|\Delta\theta\|$ and $\|\Delta\eta\|$. When $\|\Delta\eta\|\gg\|\Delta\theta\|$, Equation 125 suggests that $-\bar{\mathfrak{g}}_{\xi}(\theta)$ may not be a descent direction for $\|\Delta\theta\|$; when $\|\Delta\theta\|\gg\|\Delta\eta\|$, Equation 116 suggests that $-\bar{\mathfrak{g}}_{\xi}(\eta)$ may not be a descent direction for $\|\Delta\eta\|$.
However, combining Equations 125 and 116 gives
@@
The above inequality suggests that $-\bar{\mathfrak{g}}_{\xi}(\xi)$ is a descent direction for $\|\Delta\xi\|$ if the Lipschitz contants are small enough. ∎
@@ F.2 Proof of Theorem 1
Finally, plugging the bounds of Equations 201, 205 and 214 back into Equations 197, 198 and 199 gives
@@
Now we choose $\tau=\left\lceil\frac{\log\alpha_{t}^{-1}}{\log\rho^{-1}}\right\rceil$. Then, $\rho^{\tau}\leq\alpha_{t}$. We require the step-size to be non-increasing, and not too samll such that
@@ F.3 Proof of Corollaries
Combining Equations 229 and 230 gives

### v2

@@ C.7 PA-LFA and grid discretization
To demonstrate the effectiveness of PA-LFA, we compare it with grid discretization on the speed control example (Section C.2). The reference MFE is calculated using a grid discretization with a granularity of $1/200$. We only apply LFA to the population measure estimate, as this is our main focus. The measure basis is chosen as
$\psi_{i}(s)=cf_{\mathcal{N}}(0)-f_{\mathcal{N}}(\tan((s-s_{i})\pi)),$ (36)
where $f_{\mathcal{N}}$ is probability density function of the normal distribution with zero mean and variation $v$, and $s_{i}$ is the center of the basis function. Specifically, we set $c=1.2$ and $v=d_{2}/2$, and evenly distribute $\left\{s_{i}\right\}_{i=1}^{d_{2}}$.
@@
We run SemiSGD with $d_{2}$ states (grid discretization) or $d_{2}$ basis functions (PA-LFA) for $T=10^{4}$ steps. Recall that PA-LFA with a $d_{2}$ dimensional feature space has comparable operation complexity to grid discretization with $d_{2}$ states. Other parameters are the same as the general setup. The final MSE (accuracy) comparison is reported in Table 1. The MSE curves and learned population distributions are shown in Figure 13 and Figure 12, respectively. As we can see, PA-LFA achieves a better accuracy than grid discretization for all $d_{2}\leq 50$. This is because by choosing an appropriate measure basis, PA-LFA generalizes better than grid discretization, as the representation capacity of $\operatorname{span}\psi$ is larger than the grid discretization and the representation space of $\psi$ is closer to the MFE population measure. This is more evident when $d_{2}$ is small. As illustrated in Figure 12, with just a few basis functions, PA-LFA can capture the general shape of the MFE population distribution much better than grid discretization.
@@
Appendix D Notation
@@
Backtracking is an analysis technique introduced to tackle rapidly changing Markov chains (Zou et al., 2019; Zhang et al., 2023; 2024b). It considers a virtual stationary Markov chain by backtracking a period $\tau$, fixing the parameter $\xi_{t-\tau}$, and then sampling the Markovian observations with the fixed parameter. By the ergodicity of stationary Markov chains (Assumption 3), the virtual trajectory rapidly converges to the stationary distribution induced by $\xi_{t-\tau}$. We denote $\widetilde{O}_{t}$ as the virtual observation tuple on this virtual trajectory at time $t$. When we consider the semi-gradients on this virtual trajectory, we write out its dependence on $\widetilde{O}_{t}$ explicitly:
@@
$\bar{\mathfrak{g}}_{\circ}=\bar{\mathfrak{g}}_{\xi_{\circ}},\quad\bar{G}_{\circ}=\bar{G}_{\xi_{\circ}},\quad\bar{\varphi}_{\circ}=\bar{\varphi}_{\xi_{\circ}},\quad\bar{\psi}_{\circ}=\bar{\psi}_{\xi_{\circ}}.$ (45)
@@
Appendix E Proof of Proposition 1
@@
$\mu^{\pi}_{M}(s^{\prime})$ $=\int_{\mathcal{S}\times\mathcal{A}}P(s^{\prime}\,|\,s,a,M)\pi(a\,|\,s)\mu^{\pi}_{M}(s)\mathrm{d}s\mathrm{d}a$ (47)
$=\int_{\mathcal{S}\times\mathcal{A}}\left<\phi(s,a),\Omega_{M}\psi(s^{\prime})\right>\pi(a\,|\,s)\mu^{\pi}_{M}(s)\mathrm{d}s\mathrm{d}a$
$=\int_{\mathcal{S}\times\mathcal{A}}\left<\psi(s^{\prime}),\Omega_{M}^{T}\phi(s,a)\right>\pi(a\,|\,s)\mu^{\pi}_{M}(s)\mathrm{d}s\mathrm{d}a$
=$\psi(s^{\prime}),\underbrace{\int_{\mathcal{S}\times\mathcal{A}}\Omega_{M}^{T}\phi(s,a)\pi(a\,|\,s)\mu_{M}^{\pi}(s)\mathrm{d}s\mathrm{d}a}_{\eta}\right>.$
Appendix F Projected MFE and stationary point
We call a value function-population measure pair a stationary point if the mean-path semi-gradient evaluated at this point is zero. Sections 5 and H show that SemiSGD converges to a stationary point. Ideally, we want the stationary point to be an MFE (Definition 1). Propositions 1 and E show that for linear MFGs, images of $\mathcal{T}$ and $\mathcal{P}$ are within the linear spans of $\phi$ and $\psi$, indicating the linear structure of the MFE. However, this does not hold for general (non-linear) MFGs, hinting the discrepancy between the stationary point and the MFE for non-linear MFGs.
@@ Appendix F Projected MFE and stationary point
where $\Pi_{\phi}$ and $\Pi_{\psi}$ are orthogonal projection operators onto the linear spans of $\phi$ and $\psi$, respectively.
@@
where $\widehat{G}_{\xi}$ is the Gram matrix of $\phi$ w.r.t. the inner product $\langle\cdot,\cdot\rangle_{\mu_{\xi}}$. Note that $\widehat{G}_{\xi}$ is different from the TD operator ${G}_{\phi}$ or $G_{\xi}$ defined in Appendix D, which is only Gram-like.
@@
Appendix G Preliminary lemmas
@@
We first prove the last inequality in the lemma. By Mitrophanov (2005, Corollary 3.1), we have
@@
Appendix H Sample complexity analysis
H.1 Key lemmas
@@
$w_{\min}(\bar{G}_{\xi_{*}})=\min_{\|\theta\|=1}\mathbb{E}_{\xi_{*}}\left[\theta^{T}\phi(s,a)(\phi(s,a)-\gamma\phi(s^{\prime},a^{\prime}))^{T}\theta\right]\eqqcolon\min_{\|\theta\|=1}\mathbb{E}_{\xi_{*}}\left[u^{2}-\gamma uu^{\prime}\right],$ (132)
@@
$w_{\min}(\bar{G}_{\xi_{*}})\geq(1-\gamma)\min_{\|\theta\|=1}\mathbb{E}_{\xi_{*}}[u^{2}]=(1-\gamma)w_{\min}\left(\mathbb{E}_{\xi_{*}}[\phi\phi^{T}]\right)=(1-\gamma)w_{\min}\left(\hat{G}_{\xi_{*}}\right)=(1-\gamma)\lambda_{\min}\left(\hat{G}_{\xi_{*}}\right),$ (134)
where $\hat{G}_{\xi_{*}}$ is the Gram matrix of the feature map $\phi$ under the steady distribution induced by $\xi_{*}$, and the last equality uses the property of normal matrices. Plugging the above derivation into Equation 131 gives
@@
It is clear that Equations 135 and 126 are controlled by both $\|\Delta\theta\|$ and $\|\Delta\eta\|$. When $\|\Delta\eta\|\gg\|\Delta\theta\|$, Equation 135 suggests that $-\bar{\mathfrak{g}}_{\xi}(\theta)$ may not be a descent direction for $\|\Delta\theta\|$; when $\|\Delta\theta\|\gg\|\Delta\eta\|$, Equation 126 suggests that $-\bar{\mathfrak{g}}_{\xi}(\eta)$ may not be a descent direction for $\|\Delta\eta\|$.
However, combining Equations 135 and 126 gives
@@
The above inequality suggests that $-\bar{\mathfrak{g}}_{\xi}(\xi)$ is a descent direction for $\|\Delta\xi\|$ if the Lipschitz constants are small enough. ∎
@@ H.2 Proof of Theorem 1
Finally, plugging the bounds of Equations 211, 215 and 224 back into Equations 207, 208 and 209 gives
@@
Now we choose $\tau=\left\lceil\frac{\log\alpha_{t}^{-1}}{\log\rho^{-1}}\right\rceil$. Then, $\rho^{\tau}\leq\alpha_{t}$. We require the step-size to be non-increasing, and not too small such that
@@ H.3 Proof of Corollaries
Combining Equations 239 and 240 gives